The orbital period of Planet A is 16 years and its average distance to Star B is 5.25 a.u. Another planet called Planet C has an average distance of 10.3 a.u. from Star B, using Kepler’s Third Law of Planetary Motion, compute the orbital period of Planet C to Star B in days

1
Expert's answer
2022-05-02T13:41:56-0400

"\frac{T_1^2}{T_2^2}=\frac{a_1^3}{a_2^3},"

"T_1=T_2\sqrt\frac{a_1^3}{a_2^3}=44"years.